OpenStudy (anonymous):

Evaluate the expression and show work 32^1/5

OpenStudy (saifoo.khan):

2

OpenStudy (anonymous):

how did you get that?

OpenStudy (mathmate):

10^(log(32)/5) or by trial and error.

OpenStudy (saifoo.khan):

\[2\times 2\times 2\times 2\times 2=32\]

OpenStudy (anonymous):

how would I do this problem

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

\[3 \times 3 \times 3 \times 3 = 81\]\[3^4=81\]

OpenStudy (saifoo.khan):

both 4 cancel out, leaving u with 3.

OpenStudy (mertsj):